Rewrite the linear equations as a function of x.
-2x+3y=11

Answer by Theo(13342)

Equation is:

-2x+3y=11

Add 2x to both sides of the equation to get:

3y = 2x + 11

Divide both sides of the equation by 3 to get:

y = (2/3)*x + (11/3)

Let y = f(x) to get:

f(x) = (2/3)*x + (11/3)

What you get is y equals a function of x which is equal to the expression (2/3)*x + (11/3).

x is the argument of the function.

x is the independent variable.

f(x) is the dependent variable.

Since y = f(x), y is also the dependent variable.